Please show your work with the formulas/calculations written out (for example; without using excel tables).

9. If you paid $750 for an investment 25 years ago that is now worth $2000, what was your rate of return?

10. If you save $100 per month, what annual rate of interest must you earn in order to accumulate $10,000 after 10 years?

11. If you borrow $250,000 for 10 years and your monthly payment is $1000, what rate of interest are you paying?

12. If you bought a bond five years ago for $900 which paid annual interest of $250 and sold it today for $950, what was your rate of return?
Answer #1

Using financial calculator
1.
PV=-750
N=25
PMT=0
FV=2000
CPT I/Y=4.001%

2.
N=10*12
PMT=-100
FV=10000
PV=0
CPT I/Y=-0.316%
Annual rate=-0.316%*12=-3.792%

3.
N=10*12
PMT=-1000
PV=250000
FV=0
CPT I/Y=-1.090%
Annual rate=-1.090%*12=-13.077%

4.
N=5
PMT=250
FV=950
PV=-900
CPT I/Y=28.411%